OpenStudy (anonymous):

i need help in geometry can anyone help?? i will reward medal and fan you :)

OpenStudy (anonymous):

The following is an incomplete paragraph proving that ∠WRS ≅ ∠VQT given the information in the figure where segment UV is parallel to segment WZ.: Segments UV and WZ are parallel with line ST intersecting both at points Q and R respectively According to the given information, segment UV is parallel to segment WZ while angles SQU and VQT are vertical angles. Angle VQT is congruent to angle SQU by the Vertical Angles Theorem. Because angles SQU and WRS are corresponding angles, they are congruent according to the Corresponding Angles Postulate. Finally, angle VQT is congruent to angle WRS by the _____________________. Which Property of Equality accurately completes the proof? Reflexive Substitution Subtraction Transitive

OpenStudy (anonymous):

please help

OpenStudy (nikato):

ok, based on that proof u hsve, u hsve this: <SQU=<VQT <SQU=<WRS so now u want to find which property will help u get <VQT=<WRS

OpenStudy (anonymous):

ok .... but i'm not exactly familiar with property of equality so....

OpenStudy (nikato):

do u understand ur choices, like what each means?

OpenStudy (anonymous):

no.... :(

OpenStudy (nikato):

ok then, i will quickly go over each

OpenStudy (anonymous):

thank you

OpenStudy (nikato):

reflexive- a=a or m=m its just saying its equal to itself

OpenStudy (nikato):

substitution- its exactly what the word means. for example, u have a=b and b+2a=7 u would sunstitute the value of a into the second equation like thsi b+2(b)=7

OpenStudy (nikato):

subtracton property of equilaity is subtracting the same number on both sides like if u have an equaiton, lets say 5+6a=15 adn u wanted to subtract 5 on the left side. but u cant just do that, u also have to subtract 5 on the other side